OpenStudy (anonymous):

what would the graph look like for this soe? 2x + 3y = –3 x – y = –4

OpenStudy (anonymous):

Two straight lines?

OpenStudy (anonymous):

i knew they were both linear but i need to know where they cross

OpenStudy (anonymous):

oh, find y. i.e.x-y=-4 =>-y=-4-x =>y=4+x substitute this in your other eqn. and solve for x

OpenStudy (anonymous):

alternatively, make y the subject of both given equations and make them equal to each other.

OpenStudy (shayaan_mustafa):

Hi starrhunter :) How are you dear :D

OpenStudy (shayaan_mustafa):

Solve both equations simultaneously with some points. You will get your intersection point then.

OpenStudy (anonymous):

i am well thank you @Shayaan_Mustafa thank you for your help

OpenStudy (anonymous):

2x+3y=-3 x-y=-4 2x-2y=-8 5y=5 y=1 x+1=-4 x=-3 x=-3 y=1

OpenStudy (anonymous):

start by doubling x-y=-4 and subtract equations

OpenStudy (anonymous):

intersection points are y=1 and x=-3 the the graph looks like this